### v1

Abstract
Current methods to extract relational triples directly make a prediction based on a possible entity pair in a raw sentence without depending on entity recognition. The task suffers from a serious semantic overlapping problem, in which several relation triples may share one or two entities in a sentence. It is weak to learn discriminative semantic features relevant to a relation triple. In this paper, based on a two-dimensional sentence representation, a bi-consolidating model is proposed to address this problem by simultaneously reinforcing the local and global semantic features relevant to a relation triple. This model consists of a local consolidation component and a global consolidation component. The first component uses a pixel difference convolution to enhance semantic information of a possible triple representation from adjacent regions and mitigate noise in neighbouring neighbours. The second component strengthens the triple representation based a channel attention and a spatial attention, which has the advantage to learn remote semantic dependencies in a sentence. They are helpful to improve the performance of both entity identification and relation type classification in relation triple extraction. After evaluated on several publish datasets, it achieves competitive performance. Analytical experiments demonstrate the effectiveness of our model for relational triple extraction and give motivation for other natural language processing tasks.
keywords:
[…]
1 Introduction
Relational triple extraction is defined as a task to identify pairs of entities and extract their relations in the form of (subject, relation, object) or (s, r, o) from unstructured text. Instead of identifying relations between predefined entity types, the relational triple extraction task tries to find semantic relationships between any entity pairs in a sentence. It has the advantage to find all possible entity relations in a sentence and avoid the cascading failure caused by predicting entity types. In recent years, the task has achieved great attention in natural language processing and is widely adopted to support other downstream tasks such as knowledge graph construction (Dong et al., 2014), information retrieval (Sinha et al., 2015), etc.
Techniques to support relational triple extraction can be roughly divided into two frameworks: pipeline framework and joint framework. In pipeline framework (Chan and Roth, 2011; Zhong and Chen, 2021), the task is implemented as two independent subtasks: entity recognition and relation extraction. The main problem is that they suffer from error propagation and ignore the interaction between two subtasks. Recent researches focus on joint methods, which extract the triple (s, r, o) in an end-to-end multi-objective framework. In addition, to avoid the error propagation, joint models also have the ability to share model parameters between different tasks. They showed great potential to support relational triple extraction (Zheng et al., 2017; Wang et al., 2020; Ren et al., 2021; Yu et al., 2020).
Even great successes have been achieved in this field, relational triple extraction is still a challenging task, because a sentence usually contains several overlapped entity pairs. The overlapping phenomenon can be divided into four categories: normal, entity pair overlapping (EPO), single entity overlapping (SEO) and subject object overlapping (SOO) (Zheng et al., 2021). The normal refers to relational triples without overlapping. EPO denotes to two triples containing the same two entities but annotated with different relation types. In SEO, two triples share a same subject or object entity. In SSO, both subject and object are simultaneously shared by two relational triples. An example of the overlapping phenomenon is shown in Figure 1.
Figure 1: An example of overlapped relational triples.
The above sentence contains three relational triples. The “Live in” relation and “Locate in” relation share the same entity “Palo Alto”, which leads to the SEP overlap. On the other hand, the sentence also contains a SOO overlap between the “Locate in” relation and “Work for” relation, where the entity “Tesla” acts as the subject and object of the two relational triples, respectively. Because overlapped relational triples share the same contextual features, it is difficult to distinguish them, especially in a sequential sentence representation, where a relational triple is usually represented by fusing two entity representations.
Recent works have a trend to transform a sentence into a two-dimensional (2D) sentence representation, e.g., the planarized sentence representation (Geng et al., 2023) or the table filling (Wang et al., 2020; Ren et al., 2021; Zhang et al., 2017). In the 2D representation, each element denotes an entity pair representation in a sentence. The 2D sentence representation unfolds a semantic plane. It is effective to represent all entity pairs in a sentence. Under this representation, the relational triple extraction is conducted as element classification. The main problem for relational triple extraction in the 2D representation is that entity semantics are spread out in the 2D representation, which is influential on the extraction performance. Before to show the motivation of our model, we first give a 2D sentence representation in Figure 1.
Figure 2: 2D Sentence Representation
Refer to caption        $(a)$ Theory Representation
Refer to caption        $(b)$ Real Representation
Figure 2 contains two 2D representations about the sentence in Figure 1, where deep colors represent the strength of entity information in an element (also known as “span”). Sub-figure (a) is a theory sentence representation. Sub-figure (b) is a real sentence representation (b), which was generated by feeding the sentence into a well trained model.
In Figure 2, the theory representation and real representation shows the same semantic patterns, which indicate four important phenomena in relational triple extraction. First, the semantic plane is usually constructed by concatenating two word representations. Adjacent elements denote to overlapped phrases which share the same contextual features. Second, two entities of a relational triple can be separated by great distances, e.g., the triple ( “Franz”, “Live in”, “Palo Alto”). Third, in the 2D representation, the semantic information of a real entity is shared by all elements in the same column or the same row, because these elements are generated from the entity representation. Fourth, a sentence often contains several relation triples which may share one or two entities. The task suffers from a serious semantic overlapping problem. Under this condition, to advance the discriminability of of deep networks, it is important to make full use of local features relevant to a named entity and to encode long-distance semantic dependences (or global features) relevant to a named entity pair.
[…]
  • 1)
    Based on a 2D sentence representation, we propose a bi-consolidating model for extracting relational triples, which has the ability to simultaneously reinforce the local and global semantic features of a relational triple. Compared with related works, we achieved competitive performance on several public evaluation datasets.
  • 2)
    Several analytical experiments are conducted to demonstrate the effectiveness of our model. The results show details of the bi-consolidating model of the 2D sentence representation. It also gives motivations for researches in the relational triple extraction and other NLP tasks.
The rest of this paper is organized as follows. Section 2 provides an introduction to related works on joint entity and relation extraction. In Section 3, we present our proposed model in detail. The effectiveness of our model for joint entity and relation extraction is demonstrated through experiments conducted in Section 4. Furthermore, in Section 5, we delve into the role and performance of the various modules within the model. Finally, Section 6 offers a conclusion and outlines future works pertaining to the proposed model.
2 Related Work
Early studies (Zelenko et al., 2003; Zhou et al., 2005; Chan and Roth, 2011) often use a pipeline framework to support the joint entity and relation extraction task. They first identify named entities in the input texts. Then, the relation between all entity pairs are verified for extracting relations between them. Pipeline approaches have two shortcomings in entity and relation extraction. Firstly, it breaks the interaction between entity recognition and relation prediction. Secondly, the relation extraction task suffers from the error propagation caused in falsely named entities.
To address these limitations, researchers have delved into joint extraction methods (Sui et al., 2023; Ning et al., 2023; Zhang et al., 2023a), which simultaneously extract entities and relations in a unified framework. This paper mainly focuses on joint methods. In the followings, we roughly divided them into three categories: tagging based methods, Seq2Seq based methods and table filling based methods. Each category is introduced as follows.
2.1 Tagging based Methods
Tagging based methods convert the process of triple extraction into a series of interdependent sequence labeling tasks. For example, Wei et al. (2020) have proposed a CasRel model, which detects all potential head entities and subsequently applies relation-specific sequence taggers to identify their corresponding tail entities. Zhan et al. (2022) presented a comprehensive approach, known as DropRel, to tag the triple boundaries and relation types via three distinct dependence relations. Furthermore, to ensure the preservation of semantic information while generating distinct vectors for a given token pair, they proposed a dropout-normalization layer. Recently, Wang et al. (2023) proposed a relation-specific sequence tagging decoder that employs a filter attention mechanism to highlight more informative features, alongside an auxiliary matrix that amalgamates information related to entity pairs. However, it is worth noting that methods based on sequence tagging often suffer from issues related to semantic loss and error propagation.
2.2 Seq2Seq based Methods
Seq2Seq based methods treat a triple as a sequence of tokens and utilize an encoder-decoder framework to produce relational triples. This process is similar to machine translation. For example, CopyRE (Zeng et al., 2018) utilized as a copy mechanism to generate the relationship between two corresponding entities. The problem is that it fails to extract the order of relational facts within a sentence. In order to address this limitation, Zeng et al. (2019) introduced a reinforcement learning to a sequence-to-sequence model. Moreover, the CopyRE model is only capable of predicting the last word of an entity. In order to overcome this constraint, CopyMTL (Zeng et al., 2020) employed as a multi-task learning framework to address the multi-token entity problem. TDEER (Li et al., 2021) proposed a novel translating decoding schema that represents the relation as a translation operation from subjects to objects. However, the Seq2Seq-based approach poses challenges in constructing a vector with rich semantic features, particularly for long sentences.
2.3 Table Filling based Methods
Table filling based methods formulate the task of triple extraction as a table filling process, in which a table is comprised of the Cartesian product of an input sentence. For example, SPTree (Miwa and Bansal, 2016) effectively captured both word sequence and dependency tree substructure information. This is accomplished through the use of stacked bidirectional tree-structured LSTM-RNNs on top of bidirectional sequential LSTM-RNNs. TF-MTRNN (Gupta et al., 2016) proposed a novel context-aware joint entity and word-level relation extraction approach through semantic composition of words. This model is capable of modeling multiple relation instances without knowing the corresponding relation arguments in a sentence. RS-TTS (Zhang et al., 2023a) predicted potential relations to avoid computational redundancy and utilizes an efficient tagging and scoring strategy for entity decoding.
In addition to the above mentioned methods, researchers have investigated other approaches. For example, Bekoulis et al. (2018) presented a Multi-head model, which formulates the relational triple extraction task as a multi-head selection problem. Li et al. (2019) implemented the task as a multi-turn question answering task, where a Multi-turn QA model was proposed. Fu et al. (2019) employed a graph convolutional network-based method known as Graphrel. Eberts and Ulges (2020) utilized a span extraction approach. Sun et al. (2021) proposed a multitask learning-based RTE model. In the field of attention, Xu et al. (2015) proposed a visual attention method to model the importance of features in the image caption task. A residual attention network (Wang et al., 2017) was proposed with a spatial attention mechanism using downsampling and upsampling. Besides, a channel attention mechanism was proposed by SENet (Hu et al., 2018).
3 Methodology
The architecture of our model is demonstrated in Figure 3. It comprises three modules: Sentence Encoding Module, Bi-consolidating Module and Triple Generating Module.
Figure 3: Model Architecture. In the convolution block, the rounded rectangles represent the optional convolution type. The dashed line indicates the residual connections of the model.
The sentence encoding module transforms a sentence into a two-dimensional representation. In the bi-consolidating module, two components are designed to reinforce the local and global semantic features, respectively. Finally, the triple generating module verifies and outputs extracted relational triples. Each module is discussed as follows.
3.1 Sentence Encoding Module
Given a sentence consisting of $N$ words, it can be denoted as $\mathbf{S}=[x_{1},x_{2},\ldots,x_{N}]$ or $\mathbf{S}=x_{[1:N]}$. To capture the semantic information of each word, a commonly adopted approach is to utilize a pretrained language model that maps each word into an abstract representation. In our experiments, each sentence is fed into the BERT (Devlin et al., 2019) for learning the high-order abstract semantic features of the words:
$=\text{BERT}\left([x_{1},x_{2},\ldots,x_{N}]\right)$ (1)
$\text{or},\ \ \mathbf{H}$ $=\text{BERT}\left(\mathbf{S}\right)$
where $\mathbf{H}\in\mathbb{R}^{N\times D_{h}}$ is a sequence of token representations. The dimension of $h_{i}$ is denoted as $D_{h}$.
We adopt a self-cross embedding (Geng et al., 2023) to map the sequence $\mathbf{H}$ into a 2D sentence representation, represented as a matrix $\mathbf{M}^{B}$. It is represented as:
$\mathbf{M}^{B}=\emph{Crossing}(\mathbf{H},\mathbf{H})$ (2)
where, $\mathbf{M}^{B}_{ij}=h_{i}\oplus h_{j}$. The symbol $\oplus$” denotes the concatenation operation. The element $\mathbf{M}^{B}_{ij}$ is a span representation of $[x_{i},\ldots,x_{j}]$ (or $x_{[i:j]}$) in $\mathbf{S}$, denoted as a vector with $2D_{h}$ dimensions. Therefore, $\mathbf{M}^{B}$ is a $\mathbb{R}^{N\times N\times 2D_{h}}$ matrix.
To learn structural features of a sentence, a position embedding and an attention embedding are introduced to generate a position matrix($\mathbf{M}^{P}$) and an attention matrix ($\mathbf{M}^{A}$). The purpose of position embedding is to capture the positional information or order of tokens within a sentence. On the other hand, attention embedding is derived from the raw input using an attention layer. This embedding is learned by assigning weights to input features based on their relevance or importance to the output. The two embedding are represented as follows:
[…]
where $P_{ij}=N-i$ if $j\succ i$, else $P_{ij}=j-N$. Therefore, $\mathbf{M}^{P}$ and $\mathbf{M}^{A}$ are a $\mathbb{R}^{N\times N\times D_{p}}$ matrix and $\mathbb{R}^{N\times N\times D_{a}}$ matrix, respectively.
Finally, the 2D sentence representation is generated as:
$\mathbf{M}^{so}=\sigma(W_{so}(\mathbf{M}^{P}\oplus\mathbf{M}^{A}\oplus\mathbf{M}^{B})+b_{so})$ (4)
where $W_{so}\in\mathbb{R}^{(D_{p}+D_{a}+2D_{h})\times D_{h}}$ and $b_{so}\in\mathbb{R}^{D_{h}}$ are the trainable parameters. $\sigma$ is the ELU activation function.
3.2 Bi-consolidating Module
The quality of entity representations and relation representations are influential on the performance of the relational triple extraction. Entity representations heavily depend on the local contextual features present in a sentence. On the other hand, relations encompass semantic expressions that pertain to the entirety of a sentence. Therefore, the global features are more important for relation extraction.
To utilize local and global features, we design a novel architecture to advance the discriminability of a neural network, which is composed of a local consolidation component and a global consolidation component. Specifically, a pixel difference convolution is introduced to enhance the local information, which is implemented in adjacent regions of neighbours, thereby reinforcing local semantic features. To strengthen global semantic features, we incorporate a channel attention and a spatial attention to capture remote semantic dependencies. The local consolidation and global consolidation are discussed as follows.
3.2.1 Local Consolidation
The local consolidation module consists of a stack of multiple convolution blocks. A convolution block comprises two components: pixel difference convolution and layer normalization (Ba et al., 2016). The convolution module is constructed by stacking ${L}$ identical convolution blocks. The default value of ${L}$ is four in our experiments. The pixel difference convolution and layer normalization layer are represented as follows:
Pixel Difference Convolution: the process of pixel difference convolution (PDC) is represented as Figure 5, where the convolution kernels operate on the local feature map patch and replace the original pixels with pixel differences.
$u=f(\nabla\boldsymbol{z},\boldsymbol{\theta})=\sum_{\left(z_{i},z_{i}^{\prime}\right)\in\mathcal{P}}w_{i}\cdot\left(z_{i}-z_{i}^{\prime}\right)$ (5)
where, $z_{i}$ and $z_{i}^{\prime}$ are the input pixels, which denote to elements in the 2D sentence representation ($\mathbf{M}^{so}$). $w_{i}$ is the weight of the $k\times k$ convolution kernel. $\mathcal{P}=$$\{(z_{1},z_{1}^{\prime}),(z_{2},z_{2}^{\prime})$ $,\ldots,(z_{m},z_{m}^{\prime})\}$ ($m\leq k\times k$) is the set of pixel pairs picked from a local patch in $\mathbf{M}^{so}$.
Various strategies can be employed to select pixel pairs from the representation $\mathbf{M}^{so}$. In our study, we adopt three kinds of pixel difference convolution in Su et al. (2021), named as: central PDC (CPDC), angular PDC (APDC) and radial PDC (RPDC). We also propose five new strategies to encode pixel relations from different directions (XY-axis, diagonal, omnidirectional and bidirectional) to improve the ability for capturing the local dependencies between neighbouring elements. The strategies to select pixel pairs are as illustrated in Figure 4, in which each patch denotes to a convolution kernel. An arrow between two elements represent a pixel pair from $(z_{i},z_{i}^{\prime})$.
Figure 4: From expanded LBP descriptors, eight examples of pixel difference convolution were created.
As shown in Figure 4, in the first row, the CPDC-based methods focus on extracting features from central and adjacent spatial positions. They are enhanced through the use of three coding directions: CPDC-XY for the XY axis, CPDC-DG for diagonal directions, and CPDC-OMNI for omnidirectional directions. In the second row, the APDC-based methods leverage semantic contextual information from neighboring elements to aid in entity boundary detection. They have two bidirectional coding directions: APDC-CW for the clockwise and APDC-CCW for the anticlockwise. In the third row, the RPDC-based methods extract informative features from the peripheral regions. They are also accompanied by three coding directions: RPDC-XY for the XY axis, RPDC-DG for diagonal directions, and RPDC-OMNI for omnidirectional directions.
To explain the notion of pixel pairs in each local patch, we introduce the APDC-CW patch as an example. As showing in the $1$-th column of the $2$-th row, the APDC-CW utilizes a kernel size of $3\times 3$. A total of 8 pixel pairs is constructed in the angular clockwise direction within the local patch. Using Equation (5), the pixel differences are derived from these paired elements, then fed into a convolution layer, which entails element-wise multiplication with kernel weights followed by the summation.
Layer Normalization: when the number of convolution blocks increases during the optimization process, the gradient descent method will change the distribution of input features. To ensure the stability of these feature distributions, a layer normalization layer is employed to normalize the output of the pixel difference convolution. This normalization process enables a larger learning rate and accelerates the convergence of the model. Moreover, it is helpful to mitigate the overfitting problem, making the training process smoother.
3.2.2 Global Consolidation
The global consolidation component is introduced in this section, aiming to facilitate the model’s capture of global information from a 2D sentence representation.
As shown in Figure 3, the global consolidation component comprises a channel attention module and a spatial attention module. Given an input sentence, the channel and spatial attentions carry out complementary attention, highlighting “what” and “where”, respectively. The channel attention is applied globally, whereas the spatial attention operates on a local scale.
Channel Attention Module: we first provide a mathematical description of the inverse 2D Discrete Cosine Transform (DCT). It is expressed as follows:
$\mathrm{A}_{\mathrm{i},\mathrm{j}}^{2d}$ $=\sum_{\mathrm{h}=0}^{\mathrm{H}-1}\sum_{\mathrm{w}=0}^{\mathrm{W}-1}\mathrm{f}_{\mathrm{h},\mathrm{w}}^{2d}\underbrace{\cos\left(\frac{\pi\mathrm{h}}{\mathrm{H}}\left(\mathrm{i}+\frac{1}{2}\right)\right)\cos\left(\frac{\pi\mathrm{w}}{\mathrm{W}}\left(\mathrm{j}+\frac{1}{2}\right)\right)}_{\text{DCT weights}}$ (6)
$=\mathrm{f}_{0,0}^{2d}\mathrm{~{}B}_{0,0}^{\mathrm{i},\mathrm{j}}+\mathrm{f}_{0,1}^{2d}\mathrm{~{}B}_{0,1}^{\mathrm{i},\mathrm{j}}+\cdots+\mathrm{f}_{\mathrm{H}-1,\mathrm{~{}W}-1}^{2d}\mathrm{~{}B}_{\mathrm{H}-1,\mathrm{~{}W}-1}^{\mathrm{i},\mathrm{j}}$
$=\underbrace{\operatorname{GAP}(\mathrm{X})\mathrm{HW}\mathrm{B}_{0,0}^{\mathrm{i},\mathrm{j}}}_{\text{utilized }}+\underbrace{\mathrm{f}_{0,1}^{2d}\mathrm{~{}B}_{0,1}^{\mathrm{i},\mathrm{j}}+\cdots+\mathrm{f}_{\mathrm{H}-1,\mathrm{~{}W}-1}^{2d}\mathrm{~{}B}_{\mathrm{H}-1,\mathrm{~{}W}-1}^{\mathrm{i},\mathrm{j}}}_{\text{discarded }}$
[…]
where DCT weights are abbreviated as $\mathrm{B}^{\mathrm{i},\mathrm{j}}_{\mathrm{h},\mathrm{w}}$. $\mathrm{f}^{2d}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}}$ is the 2D DCT frequency spectrum. $\mathrm{H}$ is the height of $\mathrm{A}^{2d}$. $\mathrm{W}$ is the width of $\mathrm{A}^{2d}$.
It can be concluded from the above equation that the previous application of channel attention solely focused on the lowest frequency component of the first term GAP (Global Average Pooling). It ignores the subsequent components expressed in Equation 6, resulting in disregarded information. However, the excessive utilization of frequency components leads to the generation of superfluous information. It is imperative to take into account the optimal selection of frequency components. Therefore, in our work, we adopt an average and a max pooling operation to improve the performance.
The averaging pooling serves to eliminate redundant information while concurrently retaining the most significant information within a sentence to the greatest degree possible. In comparison, max pooling enables the identification and focuses on the most salient features presented within a given sentence. Spatial information from feature maps is aggregated through the average and max pooling operations. It generates two distinct spatial context descriptors: $\mathbf{U}^{\mathbf{c}}_{\mathbf{avg}}$, representing the average-pooled features, and $\mathbf{U}^{\mathbf{c}}_{\mathbf{max}}$, representing the max-pooled features.
Subsequently, these descriptors ($\mathbf{U}^{\mathbf{c}}_{\mathbf{avg}}$ and $\mathbf{U}^{\mathbf{c}}_{\mathbf{max}}$) are fed into a shared feedforward neural network (FFN), followed by the utilization of a sigmoid function to generate channel attention weights $\mathbf{Q}_{\mathbf{c}}\in\mathbb{R}^{C\times 1\times 1}$. The channel attention is computed as:
$\mathbf{Q}_{\mathbf{c}}$ $=\sigma\left(\emph{FFN}\left(\emph{AvgPool}(\mathbf{U})\right)+\emph{FFN}\left(\emph{MaxPool}(\mathbf{U})\right)\right)$ (7)
$=\sigma\left(\mathbf{U}_{\mathbf{avg}}^{\mathbf{c}}\mathbf{W_{c}}+\mathbf{U}_{\mathbf{max}}^{\mathbf{c}}\mathbf{W_{c}}\right)$
where $\sigma$ denotes the sigmoid function, $\mathbf{W_{c}}\in\mathbb{R}^{C\times C}$. Note that the FFN weights $\mathbf{W_{c}}$ is shared for both inputs.
The shared feedforward neural network is designed to capture the cross-channel interaction, which distinguishes them from the traditional fully-connected (FC) layers with dimensionality reduction. Our ablation studies (Section 4.3) demonstrate that dimensionality reduction results in adverse effects on channel attention prediction. Therefore, reducing the number of parameters is unnecessary.
Spatial Attention Module: in contrast to channel attention focuses on learning across channels to emphasize relevant information, spatial attention involves learning across spatial dimensions to highlight important regions within an input. These two mechanisms provide a comprehensive approach to capturing the significance of different aspects within the data. Similar to channel attention, we aggregate channel information from feature maps through the utilization of two pooling operations, resulting in the generation of two 2D maps: $\mathbf{U}^{\mathbf{s}}_{\mathbf{max}}$ and $\mathbf{U}^{\mathbf{s}}_{\mathbf{avg}}$. Each map represents the average and max pooling features of each channel, respectively. These maps are then concatenated and subjected to convolution through a convolution configured as “CNN-2D”, thereby generating our 2D spatial attention map $\mathbf{Q}_{\mathbf{s}}\in\mathbb{R}^{H\times W}$. The spatial attention is computed as:
$\mathbf{Q}_{\mathbf{s}}$ $=\sigma\left({F}_{s}([\emph{AvgPool}(\mathbf{U})\oplus\emph{MaxPool}(\mathbf{U})])\right)$ (8)
$=\sigma\left({F}_{s}\left(\left[\mathbf{U}_{\mathbf{avg}}^{\mathbf{s}}\oplus\mathbf{U}_{\mathbf{max}}^{\mathbf{s}}\right]\right)\right)$
where $\sigma$ represents the sigmoid function and ${F}_{s}$ denotes a convolution operation using a filter size of $7\times 7$.
The representation ($\mathbf{M}^{so}_{i,j}$) is fed into the local consolidation component and the global consolidation component to learn local and global semantic features. The process is represented as:
$\mathbf{T}^{so}_{i,j}=\operatorname{Attention}\left(\left[\operatorname{LN}(\operatorname{Conv^{\varepsilon}}(\mathbf{M}^{so}_{i,j}))\right]_{\times 4}\right)$ (9)
[…]
3.3 Triple Generating Module
In this module, the 2D sentence representation $\mathbf{T}^{so}_{i,j}$ is fed into a two-layer perceptron to learn a table feature for each relation candidate. The table feature for a relation $r$ is denoted as $\mathbf{T}^{r}$. Each item in $\mathbf{T}^{r}$ represents the label feature for a token pair. Specifically, for a token pair $(x_{i},x_{j})$, we represent its label feature as $\mathbf{T}^{r}(i,j)$, which is computed by Equation (10).
$\mathbf{T}^{r}(i,j)=W_{r}\operatorname{ReLU}(drop(\mathbf{T}^{so}_{i,j}W+b))+b_{r}$ (10)
where $W\in\mathbb{R}^{D_{h}\times d}$, $b$ are trainable weight and bias. $W_{r}\in\mathbb{R}^{d\times 4K}$, $b_{r}$ are trainable weight and bias. Here, $K$ signifies the number of relations, 4 signifies the number of classification tags. The function $drop(\cdot)$ denotes the dropout strategy (Srivastava et al., 2014) utilized for mitigating overfitting.
Next, by utilizing the table features $\emph{TF}_{r}$, all triples can be generated. Specifically, for every relation, the method outlined in Eq. 11 is employed to populate its corresponding table.
$\hat{\operatorname{table}_{r}}(i,j)=\operatorname{softmax}\left(\emph{TF}_{r}(i,j)\right)$ (11)
${\operatorname{table}_{r}}(i,j)=\underset{y\in Y}{\operatorname{argmax}}\left(\hat{\operatorname{table}_{r}}(i,j)[y]\right)$
[…]
Finally, the filled tables are decoded. All triples are deduced using the following method. The subject can be obtained through splicing from ‘B-E’ to ‘E-E’. The object can be obtained through splicing from ‘B-B’ to ‘B-E’. They result in the acquisition of the relationship triplet. It is noteworthy that the two paired entities share a common ‘B-E’.
We define the model the objective function as follows:
$\mathcal{L}_{\text{triple }}=$ $-\frac{1}{N\times K\times N}\times$ (12)
$\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{j=1}^{N}\log\left(\hat{\operatorname{table}_{r}}(i,j)=\hat{\operatorname{gold}_{r}}(i,j)\right)$
where $\hat{\operatorname{gold}_{r}}(i,j)$ denotes the gold tag obtained from annotations.
4 Experiment
4.1 Experimental Settings
Table 1: Statistics of the datasets, where $I$ denotes the number of triples within a sentence. Note that a sentence may simultaneously hold patterns of EPO, SEO and SOO.
Category Dataset Details of Test Set
Train Valid Test Relation Normal SEO EPO SOO I=1 I=2 I=3 I=4 I>5 Triples
$\mathrm{NYT}^{*}$ 56,195 4,999 5,000 24 3,266 1,297 978 45 3,244 1,045 312 291 108 8,110
$\mathrm{WebNLG}^{*}$ 5,019 500 703 171 245 457 26 84 266 171 131 90 45 1,591
$\mathrm{NYT}$ 56,195 5,000 5,000 24 3,222 1,273 969 117 3,240 1,047 314 290 109 8,120
$\mathrm{WebNLG}$ 5,019 500 703 216 239 448 6 85 256 175 138 93 41 1,607
[…]
In these datasets, both NYT and WebNLG have two distinct versions. Specifically, $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ are annotated with the tail tokens of a given entity. $\mathrm{NYT}$ and $\mathrm{WebNLG}$ are annotated with the whole span of an entity. Table 1 presents the statistical information of these datasets.
To evaluate the ability of our model for detecting overlapped relation instances, we follow the Zeng et al. (2018) to partition the test set based on patterns of overlapped triples. Additionally, Wei et al. (2020) highlighted that NYT and WebNLG datasets are comparatively renowned than the NYT10 and NYT11 datasets. Because NYT10 and NYT11 mainly consist of relation instances belong to the Normal class. They are primarily utilized in the main experiment to show the generalization of our model.
To evaluate the performance of our models, we utilized Partial Match and Exact Match for the NYT, NYT10, NYT11 and WebNLG datasets. Partial Match considers a triple (subject, relation, object) to be correct only if the relation, as well as the tail tokens of both subject and object, is accurate. On the other hand, Exact Match deems a triple correct only if the entire span of both entities and the relation precisely match.
Evaluation Metrics: in our experiments, we follow the same measurement in previous studies (Sui et al., 2023; Ren et al., 2022; Zheng et al., 2021). The standard metrics of micro precision $P$, recall $R$, and $F1$-score are adopted to measure the performance. They are computed as follows:
$P$ $=\frac{TP}{TP+FP}\times 100\%$ (13)
$R$ $=\frac{TP}{TP+FN}\times 100\%$
$F1$ $=\frac{2\times P\times R}{P+R}\times 100\%$
where $TP$ represents the number of correctly identified positive instances. $FN$ refers to the number of falsely identified negative instances. $FP$ indicates the number of falsely identified positive instances.
Implementation Details: in order to ensure fair comparison, we utilize the Bert-base-cased model  [^1] Available at https://huggingface.co/bert-base-cased, consisting of 12 Transformer blocks with a hidden size of $D_{h}=768$. For the WebNLG dataset, a batch size of 6 is set, whereas for other datasets, the batch size is configured to be 5. We configure the local consolidation module as ‘$[C-A_{r}-R-V]$’ and optimize all parameters using the Adam (Kingma and Ba, 2014) with the learning rate of $1\times 10^{-5}$. In equation 10, a hidden layer $d=3\times D_{h}$ d is utilized for the NYT and $\mathrm{NYT}^{*}$ dataset, whereas for the remaining datasets, a hidden layer $d=2\times D_{h}$ is employed. Additionally, the maximum sentence length for input is set at 100.
4.2 Main Results
In this experiment, our model is compared with fourteen related works. In these models, GraphRel, NovalTagging, CopyRE, ETL-span, and MultiHead utilized LSTM networks as the sentence encoder. Other models applied pre-trained BERT models for generating token representations. The performance is evaluated in terms of the Partial Match and the Exact Match. The result is presented in Table 2.
Table 2: Comparing with Related Works.
Model $\mathrm{PartialMatch}$ $\mathrm{ExactMatch}$
$\mathrm{NYT}^{*}$ $\mathrm{WebNLG}^{*}$ $\mathrm{NYT}$ $\mathrm{WebNLG}$
Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1
NovalTagging (Zheng et al., 2017) - - - - - - 32.8 30.6 31.7 52.5 19.3 28.3
CopyRE (Zeng et al., 2018) 61.0 56.6 58.7 37.7 36.4 37.1 - - - - - -
MultiHead (Bekoulis et al., 2018) - - - - - - 60.7 58.6 59.6 57.5 54.1 55.7
GraphRel (Fu et al., 2019) 63.9 60.0 61.9 44.7 41.1 42.9 - - - - - -
ETL-span (Yu et al., 2020) 84.9 72.3 78.1 84.0 91.5 87.6 85.5 71.7 78.0 84.3 82.0 83.1
CasRel (Wei et al., 2020) 89.7 89.5 89.6 93.4 90.1 91.8 - - - - - -
PMEI (Sun et al., 2021) 90.5 89.8 90.1 91.0 92.9 92.0 88.4 88.9 88.7 80.8 82.8 81.8
TPLinker (Wang et al., 2020) 91.3 92.5 91.9 91.8 92.0 91.9 91.4 92.6 92.0 88.9 84.5 86.7
StereoRel (Tian et al., 2021) 92.0 92.3 92.2 91.6 92.6 92.1 92.0 92.3 92.2 - - -
SPN (Sui et al., 2023) 93.3 91.7 92.5 93.1 93.6 93.4 92.5 92.2 92.3 - - -
PRGC (Zheng et al., 2021) $\mathbf{93.3}$ 91.9 92.6 94.0 92.1 93.0 $\mathbf{93.5}$ 91.9 92.7 89.9 87.2 88.5
TDEER (Li et al., 2021) 93.0 92.1 92.5 93.8 92.4 93.1 - - - - - -
BiRTE (Ren et al., 2022) 92.2 $\mathbf{93.8}$ $\underline{93.0}$ 93.2 94.0 93.6 91.9 $\mathbf{93.7}$ 92.8 89.0 89.5 89.3
RS-TTS (Zhang et al., 2023a) 92.9 92.8 92.8 94.4 93.9 94.1 93.0 92.6 92.8 90.7 89.7 90.2
CLFM (Wang et al., 2023) 93.0 92.3 92.7 93.9 92.6 93.3 $\underline{93.3}$ 92.4 92.8 90.3 87.9 89.1
BTDM (Zhang et al., 2023b) $\underline{93.0}$ 92.5 92.7 94.1 93.5 93.8 93.1 92.4 92.7 90.9 90.1 90.5
BitCoin (He et al., 2023) 92.9 92.8 93.1 $\mathbf{94.4}$ $\underline{94.5}$ $\underline{94.4}$ 93.1 92.6 $\underline{92.8}$ $\underline{91.9}$ $\underline{90.5}$ $\underline{91.2}$
Ours 92.9 $\underline{93.3}$ $\mathbf{93.1}$ $\underline{94.1}$ $\mathbf{96.0}$ $\mathbf{95.0}$ 93.2 $\underline{92.6}$ $\mathbf{92.9}$ $\mathbf{92.2}$ $\mathbf{91.3}$ $\mathbf{91.7}$
Model $\mathrm{PartialMatch}$ $\mathrm{ExactMatch}$
[…]
Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1
CasRel (Wei et al., 2020) 77.7 68.8 73.0 50.1 58.4 53.9 76.8 68.0 72.1 49.1 56.4 52.5
StereoRel (Tian et al., 2021) 80.0 67.4 73.2 53.8 55.4 54.6 - - - - - -
PMEI (Sun et al., 2021) $\underline{79.1}$ 70.4 74.5 $\underline{55.8}$ 59.7 57.7 77.3 $\underline{69.7}$ 73.3 $\mathbf{54.9}$ 58.9 56.8
TPLinker (Wang et al., 2020) 78.9 $\mathbf{71.1}$ $\underline{74.8}$ $\mathbf{55.9}$ $\underline{60.2}$ $\underline{58.0}$ $\underline{78.5}$ 68.8 $\underline{73.4}$ 54.8 $\underline{59.3}$ $\underline{57.0}$
Ours $\mathbf{80.4}$ $\underline{70.8}$ $\mathbf{75.3}$ 54.7 $\mathbf{63.2}$ $\mathbf{58.6}$ $\mathbf{79.7}$ $\mathbf{70.3}$ $\mathbf{74.7}$ 54.4 $\mathbf{63.0}$ $\mathbf{58.4}$
As Table 2 shown, compared with other models, our model exhibits competitive performance across all datasets. In the $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ datasets, relation triples solely be annotated with the final word of a given entity. The improvement of our model suggests that the bi-consolidating model is successful in learning global features relevant to a relation instance. On the other hand, the improvement in the $\mathrm{NYT}$ and $\mathrm{WebNLG}$ also indicates that the bi-consolidating model has the ability to correctly extract the whole span of entities, which heavily depends local features of a sentence. As shown in Table 1, these datasets usually contain relation triples with the serious semantic overlapping problem. The outcome demonstrates the effectiveness of our model in supporting the relation triple extraction.
At the bottom of Table 2, the NYT10 and NYT11 mainly contain relation instances belong to the Normal class. In comparison to robust model TPLinker which is a model based on a local semantic feature that utilizes four different tags for entity and relation detection, our model improves the performance about 0.5%, 0.6%, 1.3%, 1.4% in F1 scores, respectively. The improvement indicates that our model also has a better generalization capability. Such results confirm the effectiveness of employing the bi-consolidating module to extract all relational triples, as it enhances the refinement of semantic features by incorporating both local and global semantic features.
The improvements of the bi-consolidating model are mainly based on three advantages: first, it effectively enhances the local feature information from adjacent regions and mitigate noise in neighbouring neighbours. Secondly, it assimilates relevant information from a global perspective and learn remote semantic dependencies in a sentence, thereby effectively addressing the joint extraction task. Lastly, the utilization of two-dimensional sentence representation gives more semantic information to the sentences and effectively addresses the significant issue of semantic overlapping in relational triple extraction.
4.3 Ablation Study
The purpose of this experiment is to evaluate the contributions of each component in the bi-consolidating model. Each component is removed individually to evaluate its influence on the final performance. The dataset used for this experiment is the $\mathrm{WebNLG}$. The outcome is presented in Table 3.
Table 3: Ablation studies were performed on the WebNLG dataset using the bi-consolidating model under seven distinct settings.
Model Prec. Rec. F1
$\mathbf{Ours}$ 92.2 91.3 91.7
 w/o Bi-consolidating Module 91.4 89.0 90.2
 w/o Global Consolidation 90.6 91.1 90.8
 w/o Local Consolidation 90.7 91.4 91.0
 w/o Spatial Attention 91.5 91.3 91.4
 w/o Channel Attention 91.0 91.6 91.3
 w/o MaxPooling 91.5 91.1 91.3
 w/o AvgPooling 91.6 90.6 91.1
 w/o Dimensionality Reduction 91.0 92.0 91.5
(1) w/o Bi-consolidating Module denotes the model removing the Bi-consolidating Module from our model. The outcomes demonstrate a substantial decline in performance. It achieves the worst performance in the ablation experiments. Because this model exhibits a dual deficiency, it can not only acquire insight into or extract local pertinent semantic features, but also exhibits limitations in comprehending the comprehensive sentence semantic information from a global point.
(2) w/o Global Consolidation denotes the model removing the Global Consolidation from the complete model. This model uses only local semantic features to enhance semantic information and loses the advantage of learning distant semantic dependencies in a sentence. Compared to our model, the performance decreases about 0.9% in the $F1$ score, which confirms the importance of using global semantic features. Our model achieves a significant improvement of approximately 4.1% in the F1 score compared to the TPLinker model. Because TPLinker focuses on utilizing local semantic features. Therefore, these findings demonstrate the ability of our model to capture local information.
(3) w/o Local Consolidation denotes the model, which mainly uses global semantic features to enhance the semantic information. In this setting, the F1 score drops about 0.7% F1 score. The result serves as compelling evidence for the efficacy of the Global Consolidation model. This module enhances important features and attenuate unimportant ones controlling the scale, thus rendering the extracted features more directed. Besides, the outcome demonstrates that our model outperforms “Global Consolidation” by a marginal improvement of 0.2% in terms of the F1 score. It indicates its stronger performance in leveraging global features over local features.
(4) w/o Spatial Attention denotes to the model excluding the Spatial Attention Module from Global Consolidation, which results in the inadequate highlighting of the emphasis on “where”. The F1 score experiences a 0.3% decrease in comparison to the performance of the complete model, thus affirming the significance of the Spatial Attention Module. Spatial attention module allows the model to direct its attention towards salient keywords linked to entities or relationships within a given sentence, thereby reinforcing the significance of the respective word vectors.
(5) w/o Channel Attention denotes the model removing the Channel Attention Module from Global Consolidation. Under this condition, it fails to effectively highlight feature information about “what”. Our experiments showed that the model performed slightly lower than the “w/o Spatial Attention”, which indicates that channel attention module plays a vital role in Global Consolidation. The channel attention module acts as a core component by simulating the capture of global semantic features. It also indicates that the channel attention is applied globally, whereas the spatial attention operates on a local scale.
(6) w/o MaxPooling denotes the model without the MaxPooling operation in Global Consolidation. It hinders the model’s ability to focus on the most salient features within a given sentence. Our experimental results illustrate that the absence of MaxPooling operation has an adverse impact on both Spatial Attention Module and Channel Attention Module, ultimately leading to a decline of 0.4% in the F1 score of the model. This serves as proof that GAP can only acquire sub-optimal features due to the absence of crucial information. However, MaxPooling can effectively complement its shortcomings.
(7) w/o AvgPooling denotes the model without the AvgPooling operation in Global Consolidation. Our experiments indicate that this model performs slightly worse than "w/o MaxPooling", implying that average pooling removes some redundant information while maximizing retention of crucial information within a sentence. Therefore, AvgPooling represents the most efficient method for preserving the semantic features of a sentence.
(8) w/o Dimensionality Reduction denotes the model using the traditional fully-connected (FC) layers with dimensionality reduction in the Channel Attention Module. The experimental results illustrate that dimensionality reduction has a negative impact on channel attention prediction, leading to a 0.2% decrease in the model’s F1 score. It explains that fully connected layers without dimension reduction can enhance important features and attenuate unimportant ones controlling the scale, thus rendering the extracted features more directed.
4.4 Performance on Different Overlapping Patterns
Table 4: F1-score (%) on sentences with varying overlapping patterns and varying numbers of triples. $\mathrm{~{}I}$ represents the number of triples within a sentence. The superscript §indicates that Zheng et al. (2021) reported the results.
Dataset Model Normal $\mathrm{EPO}$ SEO SOO $\mathrm{I}=1$ $\mathrm{~{}I}=2$ $\mathrm{~{}I}=3$ $\mathrm{~{}I}=4$ $\mathrm{~{}I}\geq 5$
[…]
SPN 90.8 94.1 94.0 - 90.9 93.4 94.2 95.5 90.6
PRGC $91.0$ 94.5 94.0 81.8 91.1 93.0 93.5 95.5 93.0
Ours $\mathbf{91.2}$ $\mathbf{95.0}$ $\mathbf{94.9}$ $\mathbf{92.6}$ $\mathbf{91.2}$ $\mathbf{93.7}$ $\mathbf{93.6}$ $\mathbf{96.0}$ $\mathbf{94.3}$
[…]
SPN 89.5 90.8 94.1 - 89.5 91.3 96.4 94.7 93.8
PRGC 90.4 95.9 93.6 94.6 89.9 91.6 95.0 94.8 92.8
Ours $\mathbf{91.4}$ $\mathbf{96.7}$ $\mathbf{94.9}$ $\mathbf{96.4}$ $\mathbf{91.3}$ $\mathbf{92.7}$ $\mathbf{96.4}$ $\mathbf{95.9}$ $\mathbf{95.4}$
This experiment is conducted to validate the effectiveness of the bi-consolidating model in handling overlapping patterns and varying numbers of triples. We utilize four well-established models as comparative baselines. These models are implemented on $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ datasets.
The result in Table 4 shows that the bi-consolidating model attains the highest F1-score across all these datasets. Particularly, the performance is improved considerably in SOO and $I\geq 5$ scenarios, where the overlapping problem is more serious.
The SOO scenario has two distinct characteristics. First, nested entities are difficult to identify, due to their shared contextual features within a sentence. For example, in the triple (“Percy Smith”, Family name, “Smith”), the object entity is a substring of the subject entity. Second, the difficulty arises when the subject and object contain the same words. For instance, the sentence “The ethnic group of native Americans are from the United States where the president is the leader and baked Alaska is a dish” in $\mathrm{WebNLG}^{*}$, which contains a triple (“States”, ethnicGroup, “States”). Furthermore, compared with related works, our model shows a significant improvement in $I\geq 5$, where a sentence may hold patterns of SEO, EPO, and SOO all at once. However, our model demonstrates superior performance on both datasets for SOO and $I\geq 5$. This serves as a clear evidence to show that our model effectively addresses the overlapping triplet problem and its robustness in handling these scenarios.
4.5 Performance on Different Sub-tasks
In order to further verify the performance of the bi-consolidating model on the subtasks, we performed experiments on three subtasks in relation triple extraction: $(s,o)$, $r$ and $(s,r,o)$. The $(s,o)$ task denotes to the performance to extract entity pairs. The $r$ task is the performance of relation type classification. Finally, $(s,r,o)$ represents the relation triple extraction task. The $\mathrm{NYT}^{*}$ and $\mathrm{WebNLG}^{*}$ datasets are employed in this experiment. Table 5 presents the performance results.
Table 5: Performance on sub-tasks of relation triple extraction.
Model Element $\mathrm{NYT}^{*}$ $\mathrm{WebNLG}^{*}$
Prec. Rec. F1 Prec. Rec. F1
CasRel $(s,o)$ 89.2 90.1 89.7 95.3 91.7 93.5
$r$ 96.0 93.8 94.9 96.6 91.5 94.0
$(s,r,o)$ 89.7 89.5 89.6 93.4 90.1 91.8
SPN $(s,o)$ 93.2 92.7 92.9 95.0 95.4 95.2
$r$ 96.3 95.7 96.0 95.2 95.7 95.4
$(s,r,o)$ 93.3 91.7 92.5 93.1 93.6 93.4
PRGC $(s,o)$ $\mathbf{94.0}$ 92.3 93.1 96.0 93.4 94.7
$r$ 95.3 $\mathbf{96.3}$ 95.8 92.8 96.2 94.5
$(s,r,o)$ $\mathbf{93.3}$ 91.9 92.6 94.0 92.1 93.0
Ours $(s,o)$ 93.2 $\mathbf{94.4}$ $\mathbf{93.8}$ $\mathbf{96.2}$ $\mathbf{96.6}$ $\mathbf{96.4}$
$r$ $\mathbf{96.6}$ 95.6 $\mathbf{96.1}$ $\mathbf{96.2}$ $\mathbf{97.4}$ $\mathbf{96.8}$
$(s,r,o)$ 92.9 $\mathbf{93.3}$ $\mathbf{93.1}$ $\mathbf{94.1}$ $\mathbf{96.0}$ $\mathbf{95.0}$
Previous studies (Sui et al., 2023) have shown that the joint extraction task faces two challenges: entity pair recognition and triple formation. For example, suppose a sentence containing three instances: (“Macron”, President of, “France”), (“Macron”, Live in, “France”), (“Macron”, Place of birth, “France”). The recall on relation $r$ may decrease to 0.67 if one relation is misclassified. On the other hand, the recall on entity pair $(s,o)$ may decline to 0 if one entity pair is mistakenly recognized.
The result demonstrates that the bi-consolidating model surpasses all baseline approaches on the three sub-tasks. In the $(s,o)$ task, benefit from the ability of our model to encode local features, we achieve the best performance on recognizing entity pairs. Because the performance of entity pair identification significantly affects the overall results, we also outperform all approaches on the $(s,r,o)$ task. Compared with the $r$ task which only identifies the relation type which relies heavily on global semantic features, our model also has better performance. It further enhances the performance on the $(s,r,o)$ task.
Comparing the performance between the $\mathrm{WebNLG}^{*}$ dataset and the $\mathrm{NYT}^{*}$ dataset, the number of EPO patterns in $\mathrm{WebNLG}^{*}$ is significantly lower than $\mathrm{NYT}^{*}$ (26 vs. 978). Furthermore, the dataset $\mathrm{WebNLG}^{*}$ is annotated with a greater number of relation types compared to $\mathrm{NYT}^{*}$ (171 vs. 24). Therefore, the relation triple extraction task on $\mathrm{WebNLG}^{*}$ is more challenging than on the $\mathrm{NYT}^{*}$ task. The result shows that the bi-consolidating model achieves robust performance on the two datasets. The results demonstrate the bi-consolidating model’s robust performance on both datasets.
5 Analysis
In this section, we analyze the influence of several issues on the performance. Firstly, in Section 3.2.1, several strategies have been introduced to select pixel pairs from the representation $\mathbf{M}^{so}$. The influence of these selection strategies is analyzed. Secondly, in Section 3.2.2, the global consolidation component comprises a channel attention module and a spatial attention module. They can be implemented in a series framework or parallel framework. Finally, a visualization analysis is given to show more details about our model.
5.1 Influence of Local Consolidation Architecture
The local consolidation module is stacked with several convolution blocks. A convolution block comprises a pixel difference convolution and a normalization layer. Each pixel convolution block can employ a pixel selection strategy. Therefore, the architecture of the local consolidation module can be determined by two elements: the pixel selection strategy and the number of convolution block.
In this experiment, 9 pixel selection strategies (or PDC) is referred as: ’CPDC-XY’ ( ’$C_{xy}$’ in short), ’CPDC-DG’ (’$C_{d}$’), ’CPDC-OMNI’ (’C’ ), ’APDC-CW’ (’$A$’), ‘CNN-2D” ( ’$V$’), ’APDC-CCW’ (’$A_{r}$’), ’RPDC-XY’ (’$R_{xy}$’), ’RPDC-DG’ (’$R_{d}$’), ’RPDC-OMNI’ (’$R$’). The number of convolution blocks ranges from 1 to 4. Under this setting, the architecture of the local consolidation can be represented as a set. For example, $[C_{xy}-C_{d}-A_{r}-R]$ indicates that it is a local consolidation composed of 4 convolution blocks stacked as ’CPDC-XY’ (’$C_{xy}$’), ’CPDC-DG’ (’$C_{d}$’), ’APDC-CCW’ (’$A_{r}$’) and ’RPDC-OMNI’ (’$R$’), respectively. In this experiment, the $\mathrm{WebNLG}^{*}$ dataset is adopted. The Table 6 presents the performance of various architectures, where the top-4 configurations are highlighted in bold.
Table 6: Influence of convolution architecture.
1 Convolution Block
Architecture $[C]$ $[A_{r}]$ $[R]$
$Prec./Rec./F1$ $93.9/94.8/94.3$ $93.5/95.4/94.4$ $93.7/94.5/94.1$
2 Convolution Blocks
Architecture $[C_{xy}-C_{d}]$ $[C-C]$ $[A-A_{r}]$
$Prec./Rec./F1$ $93.7/95.3/94.5$ $92.8/95.5/94.2$ $93.2/95.5/94.3$
Architecture $[A-A]$ $[R_{xy}-R_{d}]$ $[R-R]$
$Prec./Rec./F1$ $93.1/95.4/94.2$ $93.2/95.5/94.4$ $94.2/94.3/94.3$
3 Convolution Blocks
Architecture $[C-A_{r}-R]$ $\mathbf{[C-A-A_{r}]}$ $[C-C-A_{r}]$
$Prec./Rec./F1$ $93.2/95.4/94.3$ $\mathbf{93.9}/\mathbf{95.5}/\mathbf{94.7}$ $93.1/95.2/94.2$
4 Convolution Blocks
Architecture $\mathbf{[C_{xy}-C_{d}]\times 2}$ $[A-A_{r}]\times 2$ $[R_{xy}-R_{d}]\times 2$
$Prec./Rec./F1$ $\mathbf{94.0}/\mathbf{95.6}/\mathbf{94.8}$ $93.5/94.9/94.2$ $93.6/95.4/94.5$
Architecture $[C]\times 4$ $[A_{r}]\times 4$ $[R]\times 4$
Prec./Rec./F1 $93.7/95.2/94.5$ $93.3/95.6/94.5$ $93.2/95.4/94.3$
Architecture $[C_{xy}-C_{d}-A-A_{r}]$ $[C_{xy}-C_{d}-R_{xy}-R_{d}]$ $[A-A_{r}-R_{xy}-R_{d}]$
Prec./Rec./F1 $94.0/95.1/94.6$ $93.8/95.0/94.4$ $94.1/95.2/94.6$
Architecture $[C_{xy}-C_{d}-A_{r}-R]$ $[C-C-A_{r}-R]$ $[C-A-A_{r}-R]$
Prec./Rec./F1 $93.6/95.3/94.5$ $93.8/95.1/94.5$ $94.4/94.8/94.6$
Architecture $[C-A_{r}-R_{xy}-R_{d}]$ $\mathbf{[C-A_{r}-R-R]}$ $[C-A_{r}-A_{r}-R]$
Prec./Rec./F1 $93.9/95.2/94.6$ $\mathbf{93.7}/\mathbf{95.6}/\mathbf{94.7}$ $93.9/95.2/94.6$
Architecture $\mathbf{[C-A_{r}-R-V]}$ $[C_{xy}-C_{d}-R_{d}-V]$ $[C_{xy}-C_{d}-A_{r}-V]$
Prec./Rec./F1 $\mathbf{94.1}$/$\mathbf{96.0}$/$\mathbf{95.0}$ $93.7/95.4/94.5$ $93.8/95.1/94.5$
The result shows that increasing the number of convolution blocks improves the performance. Stacking identical types of PDCs provides additional performance gains without degrading the model. The reason is that multi-stacked convolution blocks have two advantages to support the task. First, multi-stacked convolutions possess the ability to learn higher-level abstract features, which are beneficial in supporting the task of relation triple extraction. Second, stacked blocks can enlarge the receipt field, making it effective for learning semantic dependencies within a sentence.
The Top-4 configurations are ‘$[C-A-A_{r}]$’, ‘$[C_{xy}-C_{d}]\times 2$’, ‘$[C-A_{r}-R-R]$’ and ‘$[C-A_{r}-R-V]$’, which utilize various types of convolutions for capturing semantic information in different encoding directions. Specifically, the stacking of three layers of ‘$[C-A-A_{r}]$’ demonstrates the superiority of bidirectional APDCs, which may leverage semantic contextual information from adjacent elements to aid in entity boundary recognition. Similarly, the utilization of four layers of ‘$[C_{xy}-C_{d}]\times 2$’ not only underscores its prioritization of extracting features from central and adjacent positions but also showcases the substantiated efficiency of acquiring information from both the XY axis and diagonal directions. The characteristic of ‘$[C-A_{r}-R-V]$’ is the stack of four omnidirectional PDCs, which exhibits the highest performance. This can be attributed to the comprehensive utilization of four distinct encoding directions of PDCs. Therefore, the ‘$[C-A_{r}-R-V]$’ configuration is employed as the default for our model.
5.2 Influence of Global Consolidation Architecture
The global consolidation component contains a channel attention module and a spatial attention module, which can be set in a series or parallel framework. To show the influence of the global consolidation architecture, in this experiment, we evaluate the model’s performance by organizing them in a series or parallel framework. In the series framework, the output of the sentence encoding module is sequentially processed by the channel attention module and spatial attention module, respectively. In the parallel framework, the channel attention and spatial attention are paralleled. Furthermore, in the whole bi-consolidating module, the local consolidation component and the global consolidation component can be placed into a series or a parallel framework too. It is also evaluated in this section. This experiment employs the $\mathrm{WebNLG^{*}}$ dataset. The influence of the architectures is listed in Table 7, where the symbol “&” signifies a parallel connection, while “$+$” is employed to denote a series connection.
Table 7: Performance with different global consolidation architecture.
Model Prec. Rec. F1
Within Global Consolidation
 Channel & Spatial in parallel $\mathbf{94.5}$ 95.3 94.9
 Channel $+$ Spatial 94.1 $\mathbf{96.0}$ $\mathbf{95.0}$
 Spatial $+$ Channel 93.5 95.8 94.6
Within Bi-consolidating Module
 Local & Global in parallel 93.7 95.7 94.7
 Local $+$ Global $\mathbf{94.1}$ $\mathbf{96.0}$ $\mathbf{95.0}$
Table 7 presents the experimental results of different methods for arranging modules, which indicates that arranging the channel attention and spatial attention in a series framework leads to better performance compared to parallel framework. The reason is that sequentially processing the input has the advantage to learn the interaction between channel and spatial semantic space. Moreover, the channel-prior order demonstrates superior performance over the spatial-prior order. This trend can be seen in the Bi-consolidating Module too, where sequentially generating a refined subject-object features also yields better results than implementing them in parallel.
5.3 Visualization Of Different Bi-consolidating Module Components
Figure 5: The visualization of semantic overlap surrounding the entities of various components of the bi-consolidating module.
A 2D sentence representation unfolds a two-dimensional semantic space. This approach effectively expresses the semantic structures of a sentence and provides additional semantic information. The 2D representation is fed into the bi-consolidating module to generate a fine-grained subject-object features, encoded with local and global semantic features. In addition to quantitative analysis, it is also informative to provide a visualized analysis to show the influence of the bi-consolidating module. In this visualization, we collect five sentences with different overlapping patterns as SEO, SOO, Normal, EPO + SOO, SEP + SOO. These sentences are collected from the $\mathrm{WebNLG^{*}}$ dataset. They are introduced as follows:
1) SEO: the sentence is “Agra airport has the location identifier VIAG and is located in Awadh, Uttar Pradesh, Bundelkhand”. It contains two relational triples as: [“Pradesh”, isPartOf, “Awadh”], [“Pradesh”, isPartOf, “Bundelkhand”]. The entity pairs (“Pradesh”, “Awadh”) and (“Pradesh”, “Bundelkhand”), with the overlapping subject “Pradesh”, belong to a single entity overlap.
2) SOO: the sentence is “A Fortress of Grey Ice is from the United States, led by President Barack Obama”. It contains two relational triple as: [“States”, leaderName, “Obama”], [“Ice”, country, “States”]. The entity pairs (“States”, “Obama”) and (“Ice”, “States”) have the entity "State" as their subject and object, respectively, resulting in a subject object overlap.
3) Normal: the sentence is “Elliot See, born July 23 1927 in Dallas, has died”. It contains a single relational triple as: [“See”, birthPlace, “Dallas”]. There is no overlap among them.
4) EPO + SOO: the sentence is “Ampara Hospital is in Eastern Province, Sri Lanka. Ranil Wickremesinghe is a leader there”. It contains three relational triples as: [“Hospital”, country, “Lanka”], [“Lanka”, leaderName, “Wickremesinghe”] and [“Hospital”, state, “Lanka”]. The entity pairs (“Hospital”, “Lanka”) and (“Lanka”, “Wickremesinghe”) have the entity "Lanka" as their subject and object, resulting in a subject object overlap.
5) SEP + SOO: the sentence is “Alan Shepard was a crew member of NASA operated Apollo 14 who died in California which is represented by Dianne Feinstein”. It contains four relational triples as: [“Shepard”, was a crew member of, “14”], [“Shepard”, deathPlace, “California”], [“California”, senators, “Feinstein”] and [“14”, operator, “NASA”]. The entity pairs (“Shepard”, “14”) and (“Shepard”, “California”) belong to a single entity overlap. The entity pairs (“Shepard”, “California”) and (“California”, “Feinstein”) have the entity "California" as their subject and object, resulting in a subject object overlap.
These sentences are first fed into the sentence encoding module to generate $\mathbf{M}^{so}\in\mathbb{R}^{N\times N\times D_{h}}$. Then, the $\mathbf{M}^{so}$ is fed into the bi-consolidating module to learn the task relevant subject-object feature representation $\mathbf{T}^{so}$. In our model, the bi-consolidating module is composed of a local consolidation component and a global consolidation component. The local consolidation contains 4 convolution blocks referred as: “CPDC-OMNI”, "APDC-CCW", "RPDC-OMNI" and "CNN-2D". To show the visualization of various components in the bi-consolidating module, we transform each 2D representation an image by implementing the Global Average Pooling operation. This resulting image can be seen as a heat map, which denotes to a semantic plane. The visualization of these semantic planes in the five sentences is shown in Figure 5.
Figure 5 is divided into five parts marked with “SEO”, “SOO”, “Normal”, “EPO + SOO” and “SEP + SOO”. Each part is composed of five images belonging to a sentence, which are further categorized into local consolidation section and global consolidation section. The local consolidation section shows the visualization of each block as "CPDC-OMNI", "APDC-CCW", "RPDC-OMNI", "CNN-2D".
It is interesting to observe that after the utilization of the convolution “CPDC-OMNI”, all entities exhibit distinct contour characteristics. For instance, in the “Normal”, the entities “See” and “Dallas” have a high intensity. Their span edges exhibit significant semantic intensity changes in the semantic plane. Next, following the convolution block "APDC-CCW", the semantic representation surrounding the entity is amplified, which guarantees the retention of all the semantic information within the sentence. Next to the convolution block "RPDC-OMNI", the semantic information within a neighborhood is intensified, which mitigates any noise that might exist in adjacent elements. Subsequently, after the convolution block "CNN-2D", the contextual semantic representation in the proximity of the entity undergoes further enhancement, simultaneously minimizing potential disturbances originating from neighboring elements.
As visualized in every part, it can be seen that all entities exhibit obvious intensity. On the other hand, non-entity regions manifest comparatively lower intensity. In the visualization of the global consolidation module, the most impressive phenomenon is that all entity semantic representations have clear contour characteristics and have a high intensity in a 2D sentence representation. This also proves that this module has the advantage to learn remote semantic dependencies in a sentence. Another interesting phenomenon is that the representation, composed of the semantic information about an entity pair, also has a concretized semantic representation.
6 Conclusion
In this paper, based on a two-dimensional sentence representation, we propose a bi-consolidating model to address the semantic overlapping problem. Our model consists of a local consolidation component and a global consolidation component. The first component uses a pixel difference convolution to enhance semantic information of a possible triple representation of adjacent regions and mitigate noise in neighbouring neighbours. The second component strengthens the triple representation based a channel attention and a spatial attention, which has the advantage to learn remote semantic dependencies in a sentence. The bi-consolidating model simultaneously reinforce the local and global semantic features relevant to each relation triple. Our experiments demonstrate that on many benchmark datasets, the bi-consolidating model exceeds the performance of the previous state-of-the-art models. In-depth analyses are carried out to comprehend the influence of the components of the bi-consolidating model on performance and to validate their significance. Future research can focus on developing this model as a generalized framework to support various related multi-objective NLP tasks.
Acknowledgements
This work is supported by National Natural Science Foundation of China under Grant No. 62066008 and No. 62066007, National Key R&D Program of China under Grant No. 2023YFC3304500 and the Major Science and Technology Projects of Guizhou Province under Grant [2024]003.